Use the binomial theorem to find the coefficient of x^a y^b in the expansion of (2x^3




− 4y^2)




7, where




a) a = 9, b = 8.




b) a = 8, b = 9.




c) a = 0, b = 14.




d) a = 12, b = 6.




e) a = 18, b = 2.




35) How many unique partitions of the word ARKANSAS are there?




36) A group of six students consists of 3 seniors, 2 juniors, and 1 sophomore. How




many unique partitions of this group of students are there by grade?

1
Expert's answer
2021-12-26T16:58:34-0500

34)

binomial theorem:

"(p+q)^n=\sum \begin{pmatrix}
 n \\
 k 
\end{pmatrix}p^{n-k}q^k"

we have:

"(2x^3-4y^2)^7"


a)

coefficient of "x^9 y^8" :

"\begin{pmatrix}
 7 \\
 4 
\end{pmatrix}(2x^3)^3(4y^2)^4=35\cdot8\cdot256x^9y^8=71680x^9y^8"


b)

coefficient of "x^8 y^9" is 0, because it is impossible to get x8 having initial term x3 (same for y)


c)

coefficient of "x^0 y^{14}" :

"\begin{pmatrix}
 7 \\
 7 
\end{pmatrix}(2x^3)^0(4y^2)^7=4^7y^{14}=16384y^{14}"


d)

coefficient of "x^{12} y^{6}" :

"\begin{pmatrix}
 7 \\
 3 
\end{pmatrix}(2x^3)^4(4y^2)^3=35\cdot16\cdot64x^{12}y^6=35840x^{12}y^6"


e)

coefficient of "x^{18} y^{2}" :

"\begin{pmatrix}
 7 \\
 1 
\end{pmatrix}(2x^3)^6(4y^2)^1=7\cdot64\cdot4x^{18}y^2=1792x^{18}y^2"


35)

a partition of a set is a grouping of its elements into non-empty subsets, in such a way that every element is included in exactly one subset.

A multinomial coefficient describes the number of possible partitions of n objects into k groups of size n1n2, …, nk.

Multinomial Coefficient = "\frac{n!}{n_1!n_2!...n_k!}"


for word ARKANSAS:

Multinomial Coefficient = "\frac{8!}{3!1!1!1!2!}=3360"

There are 3,360 unique partitions of the word ARKANSAS.


36)

Multinomial Coefficient = "\frac{6!}{3!2!1!}=60"

There are 60 unique partitions of these students by grade.